OpenStudy (anonymous):

Help me to solve this two question. 1. different ways to solve a quadratic. When is one method prefered over the other and why? What is the discriminate and its purpose? When would it yield back zero roots, one root, or no roots? 2. Explain how to graph a polynomial function with degree higher than 2. (Make sure to go in detail on each step... finding x- intercepts, y-intercepts, end behaviors.. etc..)

OpenStudy (anonymous):

Any one help me..

OpenStudy (anonymous):

1) Different methods: Factoring, Completing the Square and Quadratic Formula. Try solve the following equations by factoring. 1) x^2+4x+3=0 2) x^2+4x+4=0 3) x^2+4x+5=0 Which is/are easy for you? Which is/are not?

OpenStudy (anonymous):

number 3

OpenStudy (anonymous):

And factoring applies to numbers 1 and 2, right? Now, what can now you say about "When is factoring preferred over the other?"

OpenStudy (anonymous):

사진 찍어서 답 보내줄께요. ^^ 지금 폰 충전중. 음. 3분만. ㅎㅎ

OpenStudy (anonymous):

To: Jnulcero, for second question, the polynomial function has to be "higher than degree of 2", so it has to be at least polynomial function with degree 3. ex: f(x) =x^3 +2x + 7

OpenStudy (anonymous):

@정현돈 Sure.

OpenStudy (anonymous):

So what Jnlucero asked me was wrong??

OpenStudy (anonymous):

1st answer is right.

OpenStudy (anonymous):

2nd one, the example was wrong, but I think he know what he is doing. Even though, he is in the process of answering.

OpenStudy (anonymous):

Sorry, but I am dealing with the first question. The examples I gave above was still for the first question (to deal with the advantages of factoring over the other). :)

OpenStudy (anonymous):

ah, my mistake. I see what you are doing now, haha.

OpenStudy (anonymous):

Oh.. thats fine

OpenStudy (anonymous):

That's ok. :) Anyway, @jisuy0102 ,now you have known how factoring is advantageous over the others, try solving the previous three equations using quadratic formula. Which is/are easy for you? Which is/are not?

OpenStudy (anonymous):

its easier to do completing the square to solve it.. Is that what ur asking??

OpenStudy (anonymous):

정현돈님 혹시 저한테 파일 보내고 계신가요??

OpenStudy (anonymous):

But, actually, quadratic formula is derived from completing the square, so quadratic formula is a lot much easier than completing the square. Oh, I got it. What I am asking is that which of the three equations above, can be easily solved using quadratic formula, and which are very difficult.

OpenStudy (anonymous):

I got it! Thank u everyone

OpenStudy (anonymous):

Factoring is the easiest on most cases. On some cases, factoring is not possible. so, you can use Quadratic formula to solve the equation.

OpenStudy (anonymous):

네. ^^ 좋은 하루 지내세요~

OpenStudy (anonymous):

네~~ 안녕히 가세요(??)